you take out a 10 year student loan for 25,000 at 5.7% annual interest, compounded monthly. if you have to make monthly payments, witch finance formula would you use for this situation

1 simple intrest

2 compound intrest

3 continuous compound intrest

4 annuity with payments in

5 annuity with payments out
Answer #1

Correct answer shall be 5. Annuity with payment Out

The student loan is repaid in monthly installments and thus the amount is Computed using the formula of annuity with payment out.

The formula shall be,

Payment = Principal * Rate * \frac{(1+rate)^{no.of period}}{(1+rate)^{no. Of period}-1}

The rate shall be monthly rate (5.7/12 = 0.475%) and Period shall be (10*12=120 Months)